Outlines

Beamer Example on Parts

Till Tantau

Fakultät für Elektrotechnik und Informatik

Technical University of Berlin

April 21, 2006

Till Tantau  Beamer Example on Parts

Outlines Part I: Review of Previous Lecture Part II: Today’s Lecture

Outline of Part I

1 Summary of the Previous Lecture

Topics Learning Objectives

Outlines Part I: Review of Previous Lecture Part II: Today’s Lecture

Outline of Part II

2 The Model of Overhead-Free Computation

The Standard Model of Linear Space

3 Limitations of Overhead-Free Computation

Linear Space versus Overhead-Free Computation

Outlines Part I: Review of Previous Lecture Part II: Today’s Lecture

Outline of Part II

2 The Model of Overhead-Free Computation

The Standard Model of Linear Space

3 Limitations of Overhead-Free Computation

Linear Space versus Overhead-Free Computation

Previous Lecture

Part I

Review of Previous Lecture

Till Tantau  Beamer Example on Parts

Previous Lecture Topics Learning Objectives

This frame shows the topics treated in the last lecture.

This and that.

Previous Lecture Topics Learning Objectives

This frame shows the topics treated in the last lecture.

This and that.

Previous Lecture Topics Learning Objectives

This frame shows the last lecture’s learning objectives.

An objective. And another one.

Previous Lecture Topics Learning Objectives

This frame shows the last lecture’s learning objectives.

An objective. And another one.

Models Limitations

Part II Today’s Lecture

Models Limitations  Standard Model

2 The Model of Overhead-Free Computation

The Standard Model of Linear Space

3 Limitations of Overhead-Free Computation

Linear Space versus Overhead-Free Computation

A frame.

Models Limitations  Linear Space

2 The Model of Overhead-Free Computation

The Standard Model of Linear Space

3 Limitations of Overhead-Free Computation

Linear Space versus Overhead-Free Computation

A frame.